# BalanceKV: KV Cache Compression through Discrepancy Theory

### v1

@@ 2.3 BalanceKV: Implementing SoftmaxBalance in Streaming for Cache Compression
For a sequence of $n$ tokens and a given memory budget of $t\ll n$, we want to design a procedure which applies SoftmaxBalance to compress the exact KV cache from storing $n$ key-value embeddings to at most $t$ in the streaming setting and can compute an estimator $z_{j}$ satisfying Equation 2 for all steps $j$ in the stream. In the streaming setting one needs to consider the following aspects. As described in the previous section, one iteration of SoftmaxBalance only allows one to select a $n/2$ sized subset of $n$ key-value embeddings, which is higher than the desired budget of $t$ embeddings. This can be easily mitigated by recursively applying SoftmaxBalance $2^{\log(n/t)}$ times, each time halving the set of key-value embeddings. However, this cannot be implemented in the streaming setting as we have a limited memory budget of $t$ which prohibits us from storing all key-value embeddings during recursion.
@@
To deal with this, we use the classical merge and reduce technique used in the design of streaming algorithms [BHM+21, MG82, GLPW16]. MergeAndReduce algorithm is a recursive binary tree-based approach that allows one to implement SoftmaxBalance recursively in a streaming setting with the total memory not exceeding $O^{*}(t)$, where $O^{*}(\cdot)$ supresses polynomial in $d,\log n$ factors, under the assumption that the norms of queries and keys are bounded. The guarantees of MergeAndReduce are presented in Theorem 3.3, its pseudocode in Algorithm 3 and a visual representation in Figure 1. If the norms of all value embeddings in the stream are the same up to constant factors, that is for all $i,j\in[n]$ $0.5\leq\|v_{i}\|_{2}/\|v_{j}\|_{2}\leq 2$, then the outputs of MergeAndReduce can be used to construct an estimator $z_{j}$ satisfying our attention approximation guarantee of equation Equation 2 with precision $\varepsilon$ for $t=O^{*}(1/\varepsilon)$. However, the value embeddings may have drastically different norms.
@@
Our main algorithm BalanceKV (pseudocode in Algorithm 1) deals with this issue by grouping the key-value embeddings in the stream according to the norms of the value embeddings, running a separate instance of MergeAndReduce on each group, and combining the outputs of each instance of MergeAndReduce. BalanceKV constructs a final estimator $z_{j}$ satisfying equation Equation 2 with precision $\varepsilon$ only using $O^{*}(1/\varepsilon)$ memory and $O^{*}(1/\varepsilon^{2})$ runtime per every step $j$ of the stream, assuming the norms of query and key embeddings are bounded. Existing methods [ZHMK24] sub sample keys and values independently in the cache, and thus they have a $1/\varepsilon^{2}$ dependence on $\varepsilon$ in the total memory. The guarantees of BalanceKV are presented in Theorem 3.1.
@@ 3 Main Theoretical Results
Our main algorithm for streaming attention approximation is BalanceKV. It takes in as input a stream of $n$ tokens $(q_{1},k_{1},v_{1}),(q_{2},k_{2},v_{2}),\ldots,(q_{n},k_{n},v_{n})$ and at every step of the stream outputs an estimate $z_{j}$ to Attn$(q_{j},K_{j},V_{j})$ (see Equation 1 for the definition of Attn(.)(.)( . )) satisfying Equation 2 with precision $\varepsilon$. Assuming that the $\ell_{2}$ norms of $q_{j},k_{j}$ are at most $r$ for all $j$, BalanceKV uses total space $O^{*}(e^{2r^{2}/\sqrt{d}}\cdot 1/\varepsilon)$ and uses $O^{*}(e^{4r^{2}/\sqrt{d}}\cdot 1/\varepsilon^{2})$ runtime at each step $j$ of the stream to output $z_{j}$. Our main theorem is as follows.
@@
For any $r,\varepsilon>0$, any positive integers $n,d$, any set of tokens $(q_{1},k_{1},v_{1}),(q_{2},k_{2},v_{2}),\\ \ldots,(q_{n},k_{n},v_{n})$ where $q_{j},k_{j},v_{j}\in\mathbb{R}^{d}$ satisfy $\|q_{j}\|_{2},\|k_{j}\|_{2}\leq r$ for all $j$, consider an invocation of BalanceKV with
  • batch size $t=O^{*}\left((e^{2r^{2}/\sqrt{d}})/\varepsilon\right)$
  • compression rate $2^{-T}$ with $T=\log(n/t)$.
Then BalanceKV outputs a vector $z_{j}$ satisfying Equation 2 with probability at least $1-1/\text{poly}(n)$ at every step $j$ of the stream. It uses total memory $O^{*}\left(e^{2r^{2}/\sqrt{d}}/\varepsilon\right)$ across all steps of the stream and runtime of $O^{*}\left(e^{4r^{2}/\sqrt{d}}/\varepsilon^{2}\right)$ per step of the stream.
BalanceKV is described in Algorithm 1. At its core BalanceKV relies on our main discrepancy based algorithm, namely SoftmaxBalance– see Section 3.1 for details on SoftmaxBalance. BalanceKV uses the output of SoftmaxBalance to compute estimates of the numerator and denominator of Attn$(q_{j},K_{j},V_{j})$ and returns the desired attention approximation $z_{j}$ for each $j$– see line 16 in BalanceKV. There are two subtleties, however. First, it is important to bucket tokens in the stream according to the norm of the value vectors – see line 4. Second, a direct application of SoftmaxBalance would require too much space. To ensure small space usage, we apply a classical streaming technique, namely the MergeAndReduce algorithm on top of SoftmaxBalance to reduce the space consumption. The space reduction achieved by MergeAndReduce is by running a logarithmic number of copies of SoftmaxBalance in a tree-like fashion: see Section 3.2 for details.
To summarize, BalanceKV groups tokens in the stream according to the norms of the corresponding value embeddings, runs a separate instance of MergeAndReduce on each group, and combines the outputs of each instance to construct the final estimate for Attn$(q_{j},K_{j},V_{j})$ at each step $j\in[n]$. Next we present SoftmaxBalance and MergeAndReduce, and then give the proof of Theorem 3.1.
@@
1:input: stream of $n$ embeddings $(q_{j},k_{j},v_{j})$, diameter $r$, batch size $t$, compression rate $2^{-T}$, precision parameter $\varepsilon$.
@@
5:   /* Bucket the stream by $\|v\|_{2}$/
6:   Let $i$ be such that: $2^{i}\leq\|v_{j}\|_{2}\leq 2^{i-1}$
@@
8:   /* Erase small norm buckets */
9:   if $\|v_{j}\|_{2}\geq v_{\text{max}}$ then
10:      $v_{\text{max}}\leftarrow\|v_{j}\|_{2}$
11:   end if
@@
15:   for $l=0,\ldots,T$ do
16:      $V^{l}\leftarrow\cup_{i}C^{l}_{i}$
17:   end for
18:   Send $(k_{j},1)$ as input to MR-Denominator
@@
20:    output $z_{j}$: $z_{j}=\frac{\sum_{l=0}^{T}2^{l}\sum_{\{k,v\}\in V^{l}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v}{\sum_{l=0}^{T}2^{l}\sum_{\{k,v\}\in K^{l}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)}$
21:   $j\leftarrow j+1$
22:until token stream ends
@@ 3.1 SoftmaxBalance
We now present our main discrepancy based compression algorithm SoftmaxBalance. Given a sequence of key/value embeddings $C=\{(k_{1},v_{1}),\ldots(k_{n},v_{n})\}$ (with key and value embeddings having possibly different dimensions), the goal of SoftmaxBalance is to produce a partition of $C$ into subsets $C^{\prime},C\setminus C^{\prime}$ such that for any query $q\in\mathbb{R}^{d}$ we have that $\sum_{(k,v)\in C^{\prime}}\exp\left(\langle k,q\rangle/\sqrt{d}\right)v\approx\sum_{(k,v)\in C\setminus C^{\prime}}\exp\left(\langle k,q\rangle/\sqrt{d}\right)v$ with high probability. Without loss of generality assume that $|C^{\prime}|\leq|C|/2$; we can output $2\sum_{(k,v)\in C^{\prime}}\exp\left(\langle k,q\rangle/\sqrt{d}\right)v$ as an approximation to $\sum_{(k,v)\in C}\exp\left(\langle k,q\rangle/\sqrt{d}\right)v$, thus achieving a factor 2 compression. Its description is presented in Algorithm 2. We note that while SoftmaxBalance takes as input a sequence of key and value embeddings, it can nevertheless be used to compute the softmax normalization: we simply run it on the keys, with the corresponding value vector one-dimensional and all equal to $1$ – see line 14 of BalanceKV, where SoftmaxBalance is called within the corresponding invocation of MergeAndReduce with value vectors as all-ones vectors. The guarantees of SoftmaxBalance are as follows.
@@
9:   $\eta_{j}\leftarrow\left\{{\begin{array}[]{ll}+1&\text{with probability}\,p_{j}\\ -1&\text{otherwise}\end{array}}\right.$

### v2

@@ 2.3 BalanceKV: Implementing SoftmaxBalance in Streaming
For a sequence of $n$ tokens and a given memory budget of $t\ll n$, we aim to design a procedure which applies SoftmaxBalance to select from $n$ key-value embeddings a set of at most $t$ in the streaming setting and can compute an estimator $z_{j}$ satisfying Equation 2 for all steps $j$ in the stream. In the streaming setting one needs to consider the following aspects. As described in the previous section, one iteration of SoftmaxBalance only allows one to select a $n/2$ sized subset of $n$ key-value embeddings, which is higher than the desired budget of $t$ embeddings. This can be easily mitigated by recursively applying SoftmaxBalance $2^{\log(n/t)}$ times, each time halving the set of key-value embeddings. However, this cannot be implemented in the streaming as we have a limited memory budget of $t$ which prohibits us from storing all key-value embeddings during recursion.
To deal with this, we use the classical merge and reduce technique used in the design of streaming algorithms [BHM+21, MG82, GLPW16].
@@
MergeAndReduce algorithm is a recursive binary tree-based approach that allows one to implement SoftmaxBalance recursively in a streaming setting with the total memory not exceeding $\widetilde{O}(dt)$, where $\widetilde{O}(\cdot)$ supresses polynomial in $\log n$ factors, under the assumption that the norms of queries and keys are bounded. The guarantees of MergeAndReduce are presented in Theorem 3.3, its pseudocode in Algorithm 4 and a visual representation in Figure 2. If the norms of all value embeddings in the stream are the same up to constant factors, that is for all $i,j\in[n]$ $0.5\leq\|v_{i}\|_{2}/\|v_{j}\|_{2}\leq 2$, then the outputs of MergeAndReduce can be used to construct an estimator $z_{j}$ satisfying our attention approximation guarantee of equation Equation 2 with precision $\varepsilon$ for $t=\widetilde{O}(\sqrt{d}/\varepsilon)$. However, the value embeddings may have very different norms.
@@
Our main algorithm BalanceKV (pseudocode in Algorithm 1) deals with this issue by grouping the key-value embeddings in the stream according to the norms of the value embeddings, running a separate instance of MergeAndReduce on each group, and combining the outputs of each instance of MergeAndReduce. BalanceKV constructs a final estimator $z_{j}$ satisfying Equation 2 with precision $\varepsilon$ only using $\widetilde{O}(d\sqrt{d}/\varepsilon)$ memory and $\widetilde{O}(d^{2}/\varepsilon^{2})$ runtime per every step $j$ of the stream, assuming the norms of query and key embeddings are bounded. Existing methods [ZHMK24] subsample keys and values independently in the cache, and thus have a $1/\varepsilon^{2}$ dependence on $\varepsilon$ in total memory. The guarantees of BalanceKV are presented in Theorem 3.1.
Finally using the lower bound on the communication complexity of INDEX, we show a lower bound on the memory complexity of any algorithm for streaming attention approximation in Theorem 3.1.
@@ 3 Main Theoretical Results
Our main algorithm for streaming attention approximation is BalanceKV. It takes in as input a stream of $n$ tokens $(q_{1},k_{1},v_{1}),(q_{2},k_{2},v_{2}),\ldots,(q_{n},k_{n},v_{n})$ and at every step of the stream outputs an estimate $z_{j}$ to Attn$(q_{j},K_{j},V_{j})$ (see Equation 1 for the definition of Attn(.)(.)( . )) satisfying Equation 2 with precision $\varepsilon$. Assuming that the $\ell_{2}$ norms of $q_{j},k_{j}$ are at most $r$ for all $j$, BalanceKV uses total space $\widetilde{O}(d\sqrt{d}e^{2r^{2}/\sqrt{d}}\cdot 1/\varepsilon)$ and uses $\widetilde{O}(d^{2}e^{4r^{2}/\sqrt{d}}\cdot 1/\varepsilon^{2})$ runtime at each step $j$ of the stream to output $z_{j}$. Our main theorem is as follows.
@@
For any $r,\varepsilon>0$, any positive integers $n,d$, any set of tokens $(q_{1},k_{1},v_{1}),(q_{2},k_{2},v_{2}),\ldots,\\ (q_{n},k_{n},v_{n})$ where $q_{j},k_{j},v_{j}\in\mathbb{R}^{d}$ satisfy $\|q_{j}\|_{2},\|k_{j}\|_{2}\leq r$ for all $j$, consider an invocation of BalanceKV with
batch size $\text{batch size }t=\widetilde{O}\left(\sqrt{d}e^{2r^{2}/\sqrt{d}}/\varepsilon\right)\text{ and compression rate }2^{-T}\text{ with }T=\log(n/t).$
Then BalanceKV outputs a vector $z_{j}$ satisfying Equation 2 with probability at least $1-1/\text{poly}(n)$ at every step $j$ of the stream. It uses total memory $\widetilde{O}\left(d\sqrt{d}e^{2r^{2}/\sqrt{d}}/\varepsilon\right)$ across all steps of the stream and runtime of $\widetilde{O}\left(d^{2}e^{4r^{2}/\sqrt{d}}/\varepsilon^{2}\right)$ per step of the stream.
A pseudocode of BalanceKV is described in Algorithm 1. At its core BalanceKV relies on our main discrepancy based algorithm, namely SoftmaxBalance– see Section 3.1 for details on SoftmaxBalance. BalanceKV uses the output of SoftmaxBalance to compute estimates of the numerator and denominator of Attn$(q_{j},K_{j},V_{j})$ and returns the desired attention approximation $z_{j}$ for each streamed index $j$. There are two subtleties, however. First, it is important to bucket tokens in the stream according to the norm of the value vectors – see lines 5 and 6. Second, a direct application of SoftmaxBalance would require too much memory space. To ensure small space usage, we apply a classical streaming technique, namely the MergeAndReduce algorithm on top of SoftmaxBalance to reduce the space consumption. The space reduction achieved by MergeAndReduce is by running a logarithmic number of copies of SoftmaxBalance in a tree-like fashion. More details are introduced in Section 3.2.
To summarize, BalanceKV groups tokens in the stream according to the norms of the corresponding value embeddings, runs a separate instance of MergeAndReduce on each group, and combines the outputs of each instance to construct the final estimate for Attn$(q_{j},K_{j},V_{j})$ at each step $j\in[n]$. Next we present SoftmaxBalance and MergeAndReduce. The full proof of Theorem 3.1 is given in appendix SectionA.1. Finally we state the theorem which provides a lower bound on the memory complexity of any algorithm for streaming attention approximation below, its full proof is provided in appendix Section C.
{restatable}
theoremlowerbound Suppose that $r^{2}\leq d$. Any streaming algorithm which on input ($(\{k_{1},v_{1}\},\ldots,\{k_{n},v_{n}\},q)$, $\|q\|_{2},\|k_{i}\|_{2}\leq r$, outputs $z_{q}$ satisfying Equation 2 with probability 0.999 has space complexity
$\Omega\left(\min\left\{\frac{1}{\varepsilon^{2}},d\exp(2r^{2}/\sqrt{d})\right\}\right).$
@@
1:input: stream of $n$ tokens $(q_{j},k_{j},v_{j})$, diameter $r$, batch size $t$, compression rate $2^{-T}$, precision parameter $\varepsilon$.
@@
5:    Find an index $i$ such that $2^{i}\geq\|v_{j}\|_{2}\geq 2^{i-1}$
@@
7:   $v_{\text{max}}\leftarrow\max\left\{\|v_{j}\|_{2},v_{\text{max}}\right\}$
@@
11:    Send $(k_{j},1)$ as input to MR-Denominator
@@
13:    output: $z_{j}=\frac{\sum_{l=0}^{T}2^{l}\sum_{\{k,v\}\in V^{l}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v}{\sum_{l=0}^{T}2^{l}\sum_{\{k,v\}\in K^{l}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)}$
14:   $j\leftarrow j+1$
15:until token stream ends
@@ 3.1 SoftmaxBalance
We now present our main discrepancy based compression algorithm, SoftmaxBalance. Given a sequence of key value embeddings $C=\{(k_{1},v_{1}),\ldots(k_{n},v_{n})\}$ (with key and value embeddings having possibly different dimensions), the goal of SoftmaxBalance is to produce a partition of $C$ into subsets $C^{\prime},C\setminus C^{\prime}$ such that for any query $q\in\mathbb{R}^{d}$ we have that $\sum_{(k,v)\in C^{\prime}}\exp(\langle k,q\rangle/\sqrt{d})v\approx\sum_{(k,v)\in C\setminus C^{\prime}}\exp(\langle k,q\rangle/\sqrt{d})v$ with high probability. Without loss of generality assume that $|C^{\prime}|\leq|C|/2$, we can then output $2\sum_{(k,v)\in C^{\prime}}\exp(\langle k,q\rangle/\sqrt{d})v$ as an approximation to $\sum_{(k,v)\in C}\exp(\langle k,q\rangle/\sqrt{d})v$, thus achieving a factor 2 compression. Its description is presented in Algorithm 2 below. We note that while SoftmaxBalance takes as input a sequence of key and value embeddings, it can nevertheless be used to compute the softmax normalization: we simply run it on the keys, with the corresponding value vector one-dimensional and all equal to $1$ – see line 11 in BalanceKV, where SoftmaxBalance is called within the corresponding invocation of MergeAndReduce with value vectors as $1$s. It’s guarantees are as follows.
@@
9:   $\eta_{j}\leftarrow\left\{{\begin{array}[]{ll}+1&\text{ with probability}\,p_{j}\\ -1&\text{o.w.}\end{array}}\right.$
@@
$\left\|\sum_{\{k,v\}\in C^{\prime}}\exp\left(\frac{\langle k,q\rangle}{\sqrt{d}}\right)v-\sum_{\{k,v\}\notin C^{\prime}}\exp\left(\frac{\langle k,q\rangle}{\sqrt{d}}\right)v\right\|_{2}\leq$
$O\left(\sqrt{s}\cdot\log(ns/\delta)\cdot\exp\left(\frac{\|q\|^{2}_{2}}{2\sqrt{d}}\right)\cdot\exp\left(\max_{j\in[n]}\frac{\left\|k_{j}\right\|^{2}_{2}}{2\sqrt{d}}\right)\cdot\max_{j\in[n]}\|v_{j}\|_{2}\right).$
The proof of the above theorem uses the breakthrough result of [ALS21] for the vector balancing problem, one of the main problems in discrepancy theory. Given a set of vectors $k_{1},\ldots,k_{n}$ the result of [ALS21] produces a subset $C$ of these vectors of at most half the size such that for any vector $q$ we have that $\sum_{k\in C}\langle k,q\rangle\approx\sum_{k\in[n]\setminus C}\langle k,q\rangle$ with high probability. Our main contribution is an algorithm for the vector balancing problem with respect to the function $\exp(\langle k,\cdot\rangle/\sqrt{d})v$ as compared to $\langle k,\cdot\rangle$ in the case of [ALS21]. We defer the proof of Theorem 3.2 to Section A.2.
3.2 MergeAndReduce
As briefly mentioned above in Section 3, MergeAndReduce is a streaming version of SoftmaxBalance. The idea is to partition the stream of tokens into batches of size $t$, apply SoftmaxBalance to the batches to reduce the size of each batch by a constant factor, and then repeat recursively – see Fig. 2 in the appendix.
If we set batch size $t$ to be about $1/\varepsilon$ (see Theorem 3.3 below for the more precise setting), we obtain a streaming algorithm that approximates $\sum_{i=1}^{j}\exp(\langle k_{i},q_{j}\rangle/\sqrt{d})v_{i}$ at any point $j$ in the stream using total space $\widetilde{O}(d\sqrt{d}e^{2r^{2}/\sqrt{d}}/\varepsilon)$ and runtime $\widetilde{O}(d^{2}e^{4r^{2}/\sqrt{d}}/\varepsilon^{2})$ per step, where $r$ is an upper bound on the norms of key and query embeddings.
As before, an important aspect is that MergeAndReduce can handle value embeddings of dimension not necessarily equal to that of key and query embeddings. Thus, when run on scalars $v_{i}=1$ for all $i$, it can also be used to approximate softmax normalization at any point $j$ in the stream. This is the main subroutine used in BalanceKV to approximate Attn$(q_{j},K_{j},V_{j})$. Its pseudocode description is presented in Section A.3.1, and its proof is in Section A.3.2. The formal guarantees are
Theorem 3.3.
batch size $\text{batch size }t=\widetilde{O}(\sqrt{s}e^{2r^{2}/\sqrt{d}}/{\varepsilon})\text{ and compression rate }2^{-T}\text{ with }T=\log(n/t).$
Then MergeAndReduce on input parameters $t,r,d,s,\varepsilon$, outputs at every step $j$ of the stream subsets of key-value embedding pairs $C^{0},\ldots,C^{T}\subset C:=\{(k_{1},v_{1}),\ldots,(k_{n},v_{n})\}$ such that
$z_{j}:=\sum_{i=0}^{T}2^{i}\sum_{\{k,v\}\in C^{i}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v$
satisfies with probability at least $1-1/\text{poly}(n)$
Total memory of the algorithm is $\widetilde{O}(d\sqrt{s}e^{2r^{2}/\sqrt{d}}/{\varepsilon})$, its $j$-th iteration runtime is $\widetilde{O}(dse^{4r^{2}/\sqrt{d}}/{\varepsilon^{2}})$.
4 Experiments
In this section we now present our experimental results. The full details of all sections as well as the experimental setup and implementation can be found in Appendix B.
4.1 Ablation Studies on Single Layer Attention Approximation
We evaluate the effectiveness of BalanceKV for approximating attention in individual layers of $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$ [DJP+24] and $\mathtt{Ministral}$-$\mathtt{8B}$-$\mathtt{Instruct}$-$\mathtt{2410}$ [Mis24] on the TriviaQA dataset from LongBench [BLZ+23]. Specifically, we examine layers 1, 2, and 5 and compare against independent uniform sampling key and value embeddings.
We provide the full experimental details in Section B.1. For each layer, we approximate attention for recent tokens using a compressed cache that retains a fixed number of initial and recent embeddings, alongside intermediate ones selected via BalanceKV, and measure its relative error against exact attention. We vary the compression rate $2^{-T}\in\{1/2,1/4,1/8,1/16\}$.
As shown in Fig. 1, BalanceKV consistently yields lower relative approximation error than approximating attention by uniform sampling past key value pairs across all settings, empirically validating its advantage as predicted by Theorem 3.1.
Figure 1: Comparison of relative errors across different layers of $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$ (left) and $\mathtt{Ministral}$-$\mathtt{8B}$-$\mathtt{Instruct}$-$\mathtt{2410}$ (right) on TriviaQA dataset.
For a fixed dataset and layer, we also analyzed how the performance and runtime of BalanceKV depend on the batch size and compression rate. More precisely, we repeat the single-layer attention approximation experiment TriviaQA and layers 1 and 15 of $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$, for batch size $\in[64,128,256]$ and compression rate $2^{-T}\in[1/2,1/4,1/8]$. The results are presented in Figure 3. As this experiment suggests, the quality of attention approximation increases as the size of the block doubles, while the runtime becomes slower as also proven theoretically.
4.2 End-to-end Evalution on LongBench
Next we evaluate BalanceKV on LongBench dataset [BLZ+23], which tests long-context understanding across tasks like QA, summarization, few-shot learning, synthetic reasoning, and code completion. Specifically, we test a version of uniform length distribution (LongBench-E). During inference, we compress the key-value cache in the prefill stage using a uniform compression rate of (approximately) 0.25 across all methods, while retaining all streamed embeddings during the decoding phase. We compare against StreamingLLM [XTC+23], SnapKV [LHY+24], PyramidKV [CZG+24], and uniform sampling (see Section 4.1), using their implementations from MInference [JLZ+24]. The evaluation follows the LongBench protocol, using three pre-trained models at different scales including $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$, $\mathtt{Qwen}$-$\mathtt{2.5}$-$\mathtt{14B}$-$\mathtt{Instruct}$ and $\mathtt{Qwen}$-$\mathtt{2.5}$-$\mathtt{32B}$-$\mathtt{Instruct}$. The results are reported in Table 1.
Notably, BalanceKV consistently achieves the best overall performance among compression methods and across all models, demonstrating its effectiveness in preserving model quality for cache compression.
The full experimental setup details can be found in Section B.2.
Method qasper multi hotpotqa 2wiki gov multinews trec triviaqa samsum p.count p.ret lcc repo-p average Qwen2.5-32B-Instruct Exact (Baseline) 44.56 50.65 69.14 60.39 21.3 19.52 75.33 81.14 43.17 22.0 99.67 50.9 35.22 51.77 StreamingLLM 20.12 34.35 51.84 48.23 19.09 17.10 61.00 51.14 28.52 23.33 41.33 39.19 26.54 35.52 PyramidKV 34.47 46.33 67.78 55.92 15.16 15.39 69.33 63.24 40.36 22.67 99.33 48.32 34.35 47.13 SnapKV 36.21 46.78 66.64 57.02 16.35 16.07 70.33 77.53 41.08 22.0 99.33 49.04 35.62 48.77 Uniform 39.28 43.82 64.82 57.84 23.10 19.50 73.00 81.63 39.70 22.00 92.00 44.97 32.19 48.76 BalanceKV 40.14 43.17 64.46 58.06 22.26 20.32 73.00 80.68 41.07 22.33 92.0 44.95 32.43 48.84 Qwen2.5-14B-Instruct Exact (Baseline) 43.39 52.63 64.06 53.71 28.07 22.4 74.67 88.75 44.81 22.33 99.0 63.69 46.3 54.14 StreamingLLM 20.73 32.62 49.93 42.39 21.63 18.69 59.67 74.96 29.57 11.67 63.0 46.16 32.25 38.71 PyramidKV 31.76 46.6 62.83 50.0 19.08 17.68 65.0 85.52 42.61 22.0 99.33 60.62 44.4 49.8 SnapKV 32.95 47.53 61.96 50.34 20.29 18.28 60.67 88.75 42.97 22.67 99.33 61.32 45.84 50.22 Uniform 37.07 41.69 61.11 49.96 29.18 22.67 71.67 87.89 40.34 22.0 84.33 58.0 42.57 49.88 BalanceKV 37.02 41.96 61.74 50.9 29.26 22.64 71.67 88.1 41.14 23.67 87.67 58.64 43.63 50.62 Llama-3.1-8B-Instruct Exact (Baseline) 42.87 48.54 52.05 38.6 31.31 22.07 71.67 91.85 42.36 20.37 98.13 49.62 42.73 50.17 StreamingLLM 20.65 30.71 39.14 32.43 23.10 18.70 58.00 83.87 28.85 20.36 97.26 33.69 30.46 39.79 PyramidKV 33.86 39.75 47.12 35.96 20.03 17.78 63.67 90.76 40.21 20.4 98.97 45.25 39.51 45.64 SnapKV 33.91 42.55 49.09 36.13 20.48 17.67 62.0 91.7 40.23 20.33 98.86 46.7 39.86 46.12 Uniform 37.18 37.15 47.49 37.82 27.56 21.06 68.67 90.48 36.13 20.33 96.26 45.72 37.09 46.38 BalanceKV 35.75 37.04 46.37 36.24 27.09 20.84 69.0 90.88 37.88 20.39 96.65 48.45 41.4 46.77
Table 1: Comparison of various cache compression methods on LongBench-E using various models. The best results among compression methods for each model are highlighted in bold.
4.3 Needle-In-A-Haystack Benchmark
We evaluate BalanceKV to the “Needle-In-A-Haystack“ benchmark [Kam23], comparing it against SnapKV, PyramidKV, StreamingLLM, and uniform sampling using the $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$. The test challenges the model to retrieve a specific sentence (the “needle”) embedded at an arbitrary position within a long context (the “haystack”). Following the setup in [FPN+24], we hide the needle at varying depths, from 0% to 100% of the total context length, across documents ranging from approximately 4K to 100K tokens. As in the previous experiments, all methods are evaluated under a fixed compression ratio of approximately 0.25.
To further enhance performance, we introduce an augmented version of BalanceKV that deterministically preserves a small set of tokens whose key embeddings are strongly anti-correlated with the rest. The standard BalanceKV procedure is then applied to the remaining tokens within each layer. As a result, BalanceKV achieves an average accuracy of 0.99, outperforming SnapKV (0.83), PyramidKV (0.90), StreamingLLM (0.31), and uniform sampling (0.90). Detailed heatmaps of performance across different context lengths and needle depths are in Figure 4 in Section B.3.
4.4 System Efficiency Metrics
Finally, we measure wall-clock times for both the prefill stage (including cache compression) and the decoding stage using a random input of length $16{,}384$ tokens, followed by the generation of $1{,}024$ tokens. Results are averaged over $10$ independent runs, with the minimum runtime reported to enhance robustness. The full results are provided in Table 2 in Section B.4.
All compression methods incur some prefill overhead compared to the uncompressed baseline (Exact). While StreamingLLM achieves the fastest decoding speed, it suffers from significantly lower accuracy (see Table 1). Among the remaining methods, BalanceKV achieves the lowest prefill latency and consistently delivers the best trade-off between efficiency and accuracy. This demonstrates that our discrepancy-based approach not only scales well in theory but also brings practical gains in end-to-end system performance, making it a compelling choice for real-world deployment scenarios.
Appendix A Full proofs of the Algorithm
A.1 Proof of Theorem 3.1
Finally equipped with Theorem 3.2 and Theorem 3.3 we now state the proof of the main Theorem 3.1.
Proof of Theorem 3.1.
[…]
Performance of the algorithm. At time step $j$, procedure MR-Denominator returns subsets $K_{1},\ldots,K_{T}$ such that
$\left|\sum_{i\in[j]}\exp\left(\frac{\langle k_{i},q_{j}\rangle}{\sqrt{d}}\right)-\sum_{l=0}^{T}2^{l}\sum_{\{k,v\}\in K^{l}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)\right|\leq\varepsilon\cdot j\cdot e^{\frac{-r^{2}}{\sqrt{d}}}\leq\varepsilon\cdot\sum_{i\in[j]}\exp\left(\frac{\langle k_{i},q_{j}\rangle}{\sqrt{d}}\right).$
[…]
therefore we get
A.2 Theoretical guarantees of SoftmaxBalance
Theorem A.1 (Theorem 1.1 in [ALS21]).
[…]
2:$c\leftarrow 30\log(n/\delta)$
3:$U_{-},U_{+}\leftarrow\emptyset$
4:for $i$ from 1 and until the end of the stream do
6:      Fail
7:   end if
9:   $\varepsilon_{i}\leftarrow+$ with probability $p_{i}$, and $\varepsilon_{i}\leftarrow-$ with probability $1-p_{i}$
10:   $U_{\varepsilon_{i}}\leftarrow U_{\varepsilon_{i}}\cup\{k_{i}\}$
11:end for
12:if $|U_{+}|\leq|U_{-}|$ then
13:   output: $U_{+}$
14:else
15:   output: $U_{-}$
16:end if
Proof of Theorem 3.2.
Define for any $k\in\mathbb{R}^{d}$ an embedding function $\varphi(k):$
It is easy to see that for any two vectors $k,q\in\mathbb{R}^{d}$
and for any $k\in\mathbb{R}^{d}$
$\|\varphi(k)\|^{2}_{2}=\exp\left(\frac{\|k\|^{2}_{2}}{\sqrt{d}}\right).$
Consider the set of vectors $\varphi(k_{1})\otimes v_{1},\ldots,\varphi(k_{n})\otimes v_{n}$. Run the Self-Balancing Walk algorithm on the set of vectors $\varphi(k_{1})\otimes v_{1},\ldots,\varphi(k_{n})\otimes v_{n}$ with failure parameter set to $\delta/s$ and denote by $C^{\prime}$ and $C\backslash C^{\prime}$ the partition of $C$ returned by the algorithm. Observe that, even though vectors $\varphi(k_{i})\otimes v_{i}$ are infinite dimensional, Self-Balancing Walk still can be implemented. The algorithm never has to keep these vectors in the memory because the only operation which requires the knowledge of the embeddings  – the inner product  – can be performed if we just store vector pairs $\{k_{i},v_{i}\}$:
Denote by $e_{1},\ldots,e_{s}$ the standard orthonormal basis in $\mathbb{R}^{s}$. By Theorem A.1, for any $i\in[s]$ with probability $1-\delta/s$
and so with probability at least $1-\delta$ all of the above inequalities hold simultaneously. To simplify the right hand side, notice that $\left\|\varphi(k)\otimes v\right\|_{2}=\exp\left(\frac{\|k\|^{2}_{2}}{2\sqrt{d}}\right)\cdot\|v\|_{2}$ and $\|\varphi(q)\otimes e_{i}\|_{2}=\exp\left(\frac{\|q\|^{2}_{2}}{2\sqrt{d}}\right)$. Observe that for any $i$ $\langle\varphi(k)\otimes v,\varphi(q)\otimes e_{i}\rangle=\langle\varphi(k),\varphi(q)\rangle\cdot[v]_{i}=\exp\left(\frac{\langle k,q\rangle}{\sqrt{d}}\right)\cdot[v]_{i}$ , where by $[v]_{i}$ we denote the $i$-th coordinate of the vector $v$. Therefore, the left hand side of the expression above is simply the absolute value of the $i$-th coordinate of the vector